# Superradiant Bosons Driving Supermassive Black Hole Mergers

### v1

@@ III Dynamical Friction in Boson Cloud
$\mathrm{Cin}(z)=\int_{0}^{z}(1-\cos t)\frac{\mathrm{d}t}{t}~{}.$ (S21)
$P_{\rm DF}=4\pi\,\frac{q^{2}M_{\rm BH}^{2}}{\alpha\sqrt{(1+q)(2/x_{\ast}-r_{0}/a)}}\,\rho(x_{\ast},\theta_{\ast})\,C_{\Lambda}\!\bigl{(}k\,r_{\Lambda}\bigr{)}.$ (S22)
$kr_{\Lambda}=\begin{cases}\sqrt{x_{p}}\,\dfrac{(1-e\cos z)^{3/2}}{\sqrt{1+e\cos z}}\,,&x_{p}\leq x_{97},\\[6.0pt] \dfrac{x_{97}}{\sqrt{x_{p}}}\,\dfrac{(1-e\cos z)^{3/2}}{\sqrt{1+e\cos z}}\,,&x_{p}>x_{97},\end{cases}$ (S23)
$R_{\ast}(\varphi_{\ast})=\frac{a(1-e^{2})}{1+e\cos{(1-\zeta)\varphi_{\ast}}}~{},$ (S24)
where $\zeta$ is the precession phase angle per period. When a secondary SMBH companion is traversing the GA boson cloud, it experiences position-dependent cloud density $\rho(R_{\ast},\theta_{\ast})$. The angular position $\theta_{\ast}$ depends on the orbital phase $\varphi_{\ast}$ and inclination angle $\iota$ as $\cos{\theta_{\ast}}=\cos{\varphi_{\ast}}\sin{\iota}$.
Then, the average of Eq. (S22) over one orbital period $T$ is
$\langle P_{\rm DF}\rangle=\frac{1}{T}\int_{0}^{2\pi}P_{\rm DF}(R_{\ast}(\varphi_{\ast}),\theta_{\ast}(\varphi_{\ast},\iota))\,\frac{\mathrm{d}t}{\mathrm{d}\varphi_{\ast}}\,\mathrm{d}\varphi_{\ast}~{}.$ (S25)
Here,
$\frac{\mathrm{d}\varphi_{\ast}}{\mathrm{d}t}=\frac{\sqrt{(1+q)M_{\rm BH}a(1-e^{2})}}{R_{\ast}(\varphi_{\ast})^{2}}~{}$ (S26)
and the orbital period $T$ can be calculated as
$T=\int_{0}^{2\pi}\frac{\mathrm{d}t}{\mathrm{d}\varphi_{\ast}}\mathrm{d}\varphi_{\ast}~{}.$ (S27)
We employ the resulting Eq. (S25) in the main text to compute the orbital decay rate due to ULB clouds.
To calculate the ULB-induced dynamical friction impact on OJ287 orbital decay rate, we use the orbital parameters of OJ287 system $a=0.056\,\text{pc}$, $e=0.657$, and $\zeta=\Delta\vartheta/2\pi\simeq 0.107$ as shown in Tab. S1. Considering these parameters in Eq. (S25), we obtain $\langle P_{\rm DF}\rangle$ as a function of boson mass $\mu$ as shown in Fig. 2. Considering $\langle P_{\rm tot}\rangle-\langle P_{\rm GW}\rangle$ allows us to set novel constraints on ULB masses or identify ULB masses that can explain a potential excess as described in the main text.
Meanwhile, this boson-induced dynamical friction can speed up the orbital shrinkage in Eq. (6), which can help solve the final parsec problem. In evaluating evolution time within the final parsec, we set a constant eccentricity in orbital shrinkage, since the evolution of eccentricity in dynamical friction dominated region is highly uncertain Quinlan:1996vp ; 2022MNRAS.511.4753G . Then we integrate Eq. (6) over evolving separations to obtain its evolution time as follows,
$\tau=\int\frac{qM_{\rm BH}^{2}}{\langle P_{\mathrm{DF}}\rangle+\langle P_{\mathrm{GW}}\rangle}\frac{1}{2a^{2}}\mathrm{d}a~{}.$ (S28)
[…]
Figure S3: Allowed SMBH mass and ULB mass parameter space where binary evolution time from $0.5\,\mathrm{pc}$ to $0.1\,\mathrm{pc}$ is shorter than $\tau_{\rm pc}<1\,\mathrm{Gyr}$, alleviating the final-parsec problem.
IV SMBH binary merger rate
$\frac{\mathrm{d}^{3}n}{\mathrm{d}z\,\mathrm{d}q\,\mathrm{d}M_{\rm BH}}=\frac{\mathrm{d}^{3}n_{\rm g}}{\mathrm{d}z\,\mathrm{d}q_{\rm g}\,\mathrm{d}M_{\rm g}}\,\frac{\mathrm{d}M_{\rm g}}{\mathrm{d}M_{\rm BH}}\,\frac{\mathrm{d}q_{\rm g}}{\mathrm{d}q}\,.$ (S29)
$\frac{\mathrm{d}^{3}n_{\rm g}}{\mathrm{d}z\,\mathrm{d}q_{\rm g}\,\mathrm{d}M_{\rm g}}=\frac{\Psi(M_{\rm g},z^{\prime})}{M_{\rm g}}\,\frac{\mathcal{P}(M_{\rm g},q_{\rm g},z^{\prime})}{T_{\rm g-g}(M_{\rm g},q_{\rm g},z^{\prime})}\,\frac{\mathrm{d}t}{\mathrm{d}z^{\prime}}\,,$ (S30)
where $\Psi$ is the galaxy stellar mass function, $\mathcal{P}$ the galaxy pair fraction, and $T_{\rm g-g}$ the galaxy merger timescale. The cosmic time-redshift relation $\mathrm{d}t/\mathrm{d}z$ in a flat $\Lambda$CDM cosmology is given by
$\frac{\mathrm{d}t}{\mathrm{d}z}=\frac{1}{(1+z)\,H(z)}\,,\quad H(z)=H_{0}\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z)^{3}}\,,$ (S31)
@@ IV SMBH binary merger rate
$m_{\psi_{z}}$ $0.11$ $\beta_{t_{0}}$ $-0.5$
$s_{\psi_{0}}$ $-1.21$ $\gamma_{t_{0}}$ $-1$
$s_{\psi_{z}}$ $-0.03$ $f_{\rm g}$ $0.615$
Table S2: Parameter values used for evaluating the the SMBH binary merger rate.
$\Psi(M_{\rm g},z)$ $=\Psi_{0}\left(\frac{M_{\rm g}}{M_{\Psi}}\right)^{s_{\Psi}}\exp\left(-\frac{M_{\rm g}}{M_{\Psi}}\right)~{},$ (S32)
$\mathcal{P}(M_{\rm g},q_{\rm g},z)$ $=\mathcal{P}_{0}(1+z)^{\beta_{p_{0}}}~{},$ (S33)
$T_{\mathrm{g-g}}(M_{\rm g},q_{\rm g},z)$ $=T_{0}(1+z)^{\beta_{t_{0}}}q_{\rm g}^{\gamma_{t_{0}}}~{},$ (S34)
@@
$\log_{10}\left(\frac{\Psi_{0}}{\mathrm{Mpc}^{-3}}\right)$ $=\psi_{0}+\psi_{z}z~{},$ (S35)
$\log_{10}\left(\frac{M_{\Psi}}{M_{\odot}}\right)$ $=m_{\psi_{0}}+m_{\psi_{z}}z~{},$ (S36)
$s_{\Psi}$ $=1+s_{\psi_{0}}+s_{\psi_{z}}z~{}.$ (S37)
[…]
$\log_{10}\left(\frac{M_{\rm BH}}{M_{\odot}}\right)=8.7+1.1\log_{10}\left(\frac{M_{\mathrm{bulge}}}{10^{11}M_{\odot}}\right)~{},$ (S38)
$M_{\mathrm{bulge}}=f_{\rm g}M_{\rm g}~{}.$ (S39)
From Eqs. (S38) and (S39), the relation between the mass ratios $q_{\rm g}$ and $q$ can be found as
$q=q_{\rm g}^{1.1}~{}.$ (S40)
In Tab. S2 we display the numerical values of relevant parameters NANOGrav:2023hfp ; Alonso-Alvarez:2024gdz .
V Boson Solitons and Gravitational Waves
Figure S4: The characteristic strain of GW background produced via SMBH binaries with different impacts from ULB soliton cores and superradiant GA boson clouds. We consider GA boson clouds in $|211\rangle$ state, with critical SMBH mass satisfying $\Gamma_{211}=1/\tau_{\mathrm{U}}$ for green solid line and $\Gamma_{211}=170/\tau_{\mathrm{U}}$ for green dashed line. The shaded region depicts impact from superradiant boson clouds on SGWB, considering ULB masses in the range of $(0.9-3)\times 10^{-21}\,\text{eV}$. Impact of solitonic cores with ULB masses of $\mu=1.2,1.5\times 10^{-21}\,\mathrm{eV}$ is shown. The mass ratio $\varepsilon$ between soliton core and SMBH is set as $10^{-3}$. The observation data points come from NANOGrav NANOGrav:2023gor , EPTA EPTA:2023xxk , and PPTA Reardon:2023gzh are shown.
$\rho_{\mathrm{sol}}\simeq\rho_{0}e^{-2\alpha\mu r}$ (S41)
with
$\rho_{0}\simeq 5\times 10^{4}\varepsilon\left(\frac{M}{10^{8}\,M_{\odot}}\right)^{4}\left(\frac{\mu}{10^{-21}\,\mathrm{eV}}\right)^{6}\,M_{\odot}\,\mathrm{pc}^{-3}~{},$ (S42)
@@ V Boson Solitons and Gravitational Waves
In Fig. S4 we display the GW strain for stochastic background from SMBH solitonic cores for various ULB DM masses considering population of SMBHs following Eq. (8) and SMBH mass distribution in the range of $10^{5}-10^{10}M_{\odot}$. We compare these results with dynamical friction effects from ULB GA boson clouds in $|211\rangle$ state considering SMBHs with mass greater than critical mass $M_{\mathrm{crit}}$ that satisfies condition $\Gamma_{211}=1/\tau_{\mathrm{U}}$. The dynamical friction effects from ULB-DM solitonic cores and from superradiant clouds imprint qualitatively different low-frequency turnovers in the PTA spectrum, allowing these two scenarios to be distinguished.

### v2

@@ Appendix C SMBH binary merger rate
$\log_{10}\left(\frac{\Psi_{0}}{\mathrm{Mpc}^{-3}}\right)$ $=\psi_{0}+\psi_{z}z~,$ (48)
$\log_{10}\left(\frac{M_{\Psi}}{M_{\odot}}\right)$ $=m_{\psi_{0}}+m_{\psi_{z}}z~,$ (49)
$s_{\Psi}$ $=1+s_{\psi_{0}}+s_{\psi_{z}}z~.$ (50)
We can further relate $M_{\rm g}$ to the SMBH mass via the $M_{\rm BH}$$M_{\rm bulge}$ relation Kormendy:2013dxa
$\log_{10}\left(\frac{M_{\rm BH}}{M_{\odot}}\right)=8.7+1.1\log_{10}\left(\frac{M_{\mathrm{bulge}}}{10^{11}M_{\odot}}\right)~,$ (51)
$M_{\mathrm{bulge}}=f_{\rm g}M_{\rm g}~.$ (52)
From Eqs. (51) and (52), the relation between the mass ratios $q_{\rm g}$ and $q$ can be found as
$q=q_{\rm g}^{1.1}~.$ (53)
In Tab. 2 we display the numerical values of relevant parameters NANOGrav:2023hfp ; Alonso-Alvarez:2024gdz .
@@
$m_{\psi_{z}}$ $0.11$ $\beta_{t_{0}}$ 0.5-0.5
$s_{\psi_{0}}$ 1.21-1.21 $\gamma_{t_{0}}$ 1-1
$s_{\psi_{z}}$ 0.03-0.03 $f_{\rm g}$ $0.615$
Table 2: Parameter values used for evaluating the the SMBH binary merger rate.
Appendix D Statistical Analysis of GW Background Fit
To evaluate the performance of stochastic gravitational wave background in superradiant boson cloud scenario, we estimate the value of $\chi^{2}$ in both superradiant boson cloud scenario and pure GW scenario. The $\chi^{2}$ can be calculated as
$\chi^{2}=\sum_{i=1}^{N}\frac{(O_{i}-E_{i})^{2}}{\sigma_{i}^{2}}~,$ (54)
where $O_{i}$ is the $i$th observed characteristic strain at measured frequency, $E_{i}$ is the corresponding expected value of characteristic strain in difference scenarios, and $\sigma_{i}^{2}$ is the variance of the $i$th distribution of characteristic strain. $N$ is the total number of observed data points. Give the characteristic strain of PTA data as shown in Table. 3, we have the observed value of characteristic strain $O_{i}$ and corresponding variance $\sigma_{i}^{2}$. Since the value of upper variance $\sigma_{+}^{2}$ and lower variance $\sigma_{-}^{2}$ are different, we choose the mean value of upper and lower variance in $\chi^{2}$ calculation as
$\sigma^{2}=\frac{\sigma_{+}^{2}+\sigma_{-}^{2}}{2}~.$ (55)
Then we apply Eq. (21) to calculate $E_{i}$ at measured frequency in $O_{i}$. It gives a $\chi^{2}$ fit result from Fig. 5 that predicted strain in GA cloud scenario gives $\chi^{2}=16.5$ (blue solid - all SMBHs with maximal spin), $\chi^{2}=17.5$ (blue dashed - conservative superradiance timescales), while predicted strain in GW scenario gives $\chi^{2}=18.8$ (black solid - general relativity prediction).
$f\,[\text{yr}^{-1}]$ $h_{c}/10^{-15}$ Ref $f\,[\text{yr}^{-1}]$ $h_{c}/10^{-15}$ Ref
$0.062$ $6.5^{+4.5}_{-2.4}$ NANOGrav $0.055$ $8.3^{+6.8}_{-3.4}$ PPTA
$0.12$ $7.9^{+3.0}_{-1.8}$ $0.11$ $9.6^{+4.2}_{-3.6}$
$0.19$ $7.4^{+3.1}_{-2.0}$ $0.17$ $7.4^{+3.1}_{-1.8}$
$0.25$ $6.4^{+3.3}_{-1.9}$ $0.22$ $6.3^{+4.5}_{-2.5}$
$0.31$ $9.3^{+4.8}_{-4.2}$ $0.28$ $1.0^{+3.2}_{-0.8}$
$0.097$ $8.0^{+4.0}_{-2.7}$ EPTA $0.33$ $6.3^{+4.4}_{-4.0}$
$0.19$ $9.6^{+2.9}_{-1.9}$ $0.39$ $3.7^{+3.7}_{-1.5}$
$0.29$ $8.2^{+3.8}_{-8.2}$ $0.44$ $7.1^{+3.9}_{-2.6}$
$0.39$ $11.0^{+4.5}_{-2.9}$ $0.50$ $1.6^{+3.2}_{-1.0}$
$0.48$ $5.2^{+10.0}_{-5.5}$ $0.55$ $3.6^{+3.6}_{-2.09}$
$0.58$ $0.6^{+4.8}_{-0.4}$
Table 3: The data points of the PTA characteristic strain in NANOGrav NANOGrav:2023gor , EPTA EPTA:2023xxk and PPTA Reardon:2023gzh .
Appendix E Boson Solitons and Gravitational Waves
Figure 7: The characteristic strain of GW background produced via SMBH binaries with different impacts from ULB soliton cores and superradiant GA boson clouds. We consider GA boson clouds in $|211\rangle$ state, with critical SMBH mass satisfying $\Gamma_{211}=1/\tau_{\mathrm{U}}$ for green solid line and $\Gamma_{211}=170/\tau_{\mathrm{U}}$ for green dashed line. The shaded region depicts impact from superradiant boson clouds on SGWB, considering ULB masses in the range of $(0.9-3)\times 10^{-21}\,\text{eV}$. Impact of solitonic cores with ULB masses of $\mu=1.2,1.5\times 10^{-21}\,\mathrm{eV}$ is shown. The mass ratio $\varepsilon$ between soliton core and SMBH is set as $10^{-3}$. The observation data points come from NANOGrav NANOGrav:2023gor , EPTA EPTA:2023xxk , and PPTA Reardon:2023gzh are shown.
$\rho_{\mathrm{sol}}\simeq\rho_{0}e^{-2\alpha\mu r}$ (56)
@@ Appendix E Boson Solitons and Gravitational Waves
$\rho_{0}\simeq 5\times 10^{4}\varepsilon\left(\frac{M}{10^{8}\,M_{\odot}}\right)^{4}\left(\frac{\mu}{10^{-21}\,\mathrm{eV}}\right)^{6}\,M_{\odot}\,\mathrm{pc}^{-3}~,$ (57)
@@
In Fig. 7 we display the GW strain for stochastic background from SMBH solitonic cores for various ULB DM masses considering population of SMBHs following Eq. (21) and SMBH mass distribution in the range of $10^{5}-10^{10}M_{\odot}$. We compare these results with dynamical friction effects from ULB GA boson clouds in $|211\rangle$ state considering SMBHs with mass greater than critical mass $M_{\mathrm{crit}}$ that satisfies condition $\Gamma_{211}=1/\tau_{\mathrm{U}}$. The dynamical friction effects from ULB-DM solitonic cores and from superradiant clouds imprint qualitatively different low-frequency turnovers in the PTA spectrum, allowing these two scenarios to be distinguished.